I need help with the following problem:
The product of two consecutive positive even numbers is 624, what are the numbers?
I cannot find this anywhere in my book. Thanks so much!

Answer by mananth(16949)

Let the numbers be x & x+2
x(x+2)=624
x^2+2x =624
complete the square on the left side
x^2+2x+1= 624+1
(x+1)^2= 625
Take the square root
(x+1) = +/- 25
x= -1+25 OR -1-25
x= 24 OR -26
(24,26)
(-26,-24)
Since they are positive
the numbers are 24 & 26